# Quasi-Clique Discovery via Energy Diffusion

### v2

@@ 3 $P\leftarrow$ vertices of AA sorted by decreasing $f$;
6
7while $|S|>1$ and $\eta(G[S])<\gamma$ do
@@ 19       $\textit{changed}\leftarrow\textbf{true}$;
Let $k=|A|$ be the number of active vertices. The ExtractQuasiClique stage runs in $O(k^{2}+m(A))$, where $m(A)$ is the number of edges in the induced subgraph $G[A]$.
@@
The overall complexity is determined by the main steps of the algorithm. Sorting the $k$ active vertices takes $O(k\log k)$ time. The subsequent sweep cut, which involves calculating conductance for $k$ prefixes, can be implemented in $O(m(A))$ time with incremental updates for the cut and volume. The final refinement stage is the most computationally intensive. The pruning and expansion loops can, in the worst case, lead to a series of $O(k)$ vertex additions or removals. With incremental maintenance of internal degrees and edge counts, each step of these loops can take up to $O(k)$ time to find the best candidate to remove or to scan through remaining candidates for addition. This results in a worst-case complexity of $O(k^{2})$ for the refinement phase. Summing these costs, the total complexity is $O(k\log k+m(A)+k^{2})$, which simplifies to $O(k^{2}+m(A))$. ∎
@@
The theoretical worst-case time complexity of the EDQC algorithm (Algorithm 1) is $O(n^{3})$.
@@
The algorithm’s main loop iterates up to $n$ times. Each iteration calls two procedures: EnergyDiffusion, with a complexity of $O(T(n+m))$, and ExtractQuasiClique, with a worst-case complexity of $O(n^{2}+m)$. As $T$ is a small constant, the cost of a single iteration is dominated by the extraction step, at $O(n^{2}+m)$. Therefore, the total worst-case complexity is $O(n^{3}+nm)$. Since $m=O(n^{2})$ for any simple graph, this simplifies to $O(n^{3})$. The initial sorting of vertices by non-increasing degree is a lower-order term that does not affect this bound.
We note that this is a theoretical worst-case analysis. In practice, EDQC is run as a time-bounded heuristic, and its empirical efficiency is demonstrated in our experimental section. ∎
4 Experimental Evaluation
We conducted a comprehensive experimental evaluation to validate the performance and robustness of EDQC. This section is organized as follows. We first detail the experimental setup, the baseline methods, and the datasets used. We then present the main results on effectiveness and efficiency, followed by a statistical analysis to confirm the significance of our findings. Subsequently, we provide an in-depth analysis of the algorithm, including an ablation study to verify the contribution of its key components and an empirical study to validate its core premise. The section concludes with a summary of our parameter sensitivity analysis.
4.1 Experimental Setup
All methods were implemented in C++ and compiled with g++ 9.4.0 using the ‘-O3’ optimization flag. The experiments were conducted on a machine equipped with an AMD EPYC 7H12 CPU (1.5 GHz base frequency) and 128 GB of RAM, running Ubuntu 20.04.
Each baseline was run with the parameter settings recommended in its original paper. For EDQC, we fixed the diffusion rounds at $T=2$ and the activation threshold at $\theta=10^{-4}$. These values were selected based on a sensitivity study, the details of which are provided in Section 4.8.
Due to inherent randomness, all non-deterministic methods (all except NBSim) were executed 10 times with random seeds from 1 to 10; NBSim, being deterministic, was run once. For the non-deterministic methods, we report their average quasi-clique size and standard deviation (STD) over the 10 runs, presented in the format ‘average size $\pm$ STD’, where the standard deviation is reported only if it is non-zero. All runs were subject to a 60-second timeout. If an algorithm failed to find any valid $\gamma$-quasi-clique in all 10 runs for an instance, we mark its result as ‘N/A’. The 60-second timeout reflects common practice in recent time-bounded heuristic studies lin2017reduction ; zheng2023farsighted .
4.2 Baseline Methods
@@
  • NBSim/FastNBSim pang2024similarity : Similarity-based heuristics that expand subgraphs based on neighborhood overlap. FastNBSim accelerates NBSim via MinHash-based approximation.
  • NuQClqF1/NuQClqR1 liu2024optimization : Advanced variants of NuQClq chen2021nuqclq . These algorithms employ information feedback models that integrate historical solutions using fitness-weighted scores to guide local search.
4.3 Datasets
We evaluated all algorithms on 75 real-world datasets from SNAP snapnets and the Network Repository nr . The collection covers diverse domains, including web, social, and communication graphs. The datasets span a wide spectrum of scales, with vertex counts ($n$) ranging from 62 to over 16.7 million, and edge densities spanning from $8.41\times 10^{-2}$ down to as low as $1.87\times 10^{-7}$. This diversity provided a challenging testbed for evaluating the scalability and effectiveness of the algorithms.
4.4 Main Results
We now present the primary results comparing the effectiveness and efficiency of EDQC against the baselines. A key challenge in this comparison is that NBSim and FastNBSim do not accept a density threshold $\gamma$ as input. To facilitate a fair evaluation, we designed two separate experimental settings (Table 1 and Table 2). To save space, we omit instances where all four compared algorithms found solutions of the same average size.
Table 1: Quasi-clique sizes with the threshold $\gamma$ set by NBSim’s output.
@@ 4.4 Main Results
Table 2: Quasi-clique sizes with $\gamma$ set by FastNBSim’s max density over 10 runs. For FastNBSim, stats are on successful runs only (rate shown in parentheses).
@@
Beyond average size, EDQC demonstrates exceptional stability. It consistently reports a standard deviation close to zero (e.g., $0.32$ on the large eu-2005 graph), even on complex instances. This is in stark contrast to the NuQClq variants, which frequently exhibit high variance on the same instances (e.g., a STD of $115.84$ for NuQClqF1 on eu-2005). Furthermore, the results highlight the unreliability of the similarity-based methods. FastNBSim, for instance, often fails to achieve its own best-found density (succeeding only 1/10 times on many graphs). This lack of explicit density control makes them less suitable for applications requiring highly cohesive structures, a task where EDQC’s reliability is a distinct advantage.
Figure 2: Runtime comparison of EDQC against baselines on a log-log scale.
@@
The efficiency of EDQC is compared against baselines in Figure 2, where, following prior work sunnumds , instances where the runtime of both algorithms is below 0.1 seconds are excluded. On this log-log plot, points falling below the $1\times$, $10\times$, and $100\times$ lines indicate that EDQC is faster than the baseline by the corresponding factor or more. The results clearly show that EDQC is substantially more efficient than the high-performing local search methods, NuQClqF1 and NuQClqR1, with a large number of points falling below the $10\times$ line, signifying a speedup of over an order of magnitude. In comparison to the fast similarity-based heuristics (NBSim and FastNBSim), EDQC’s runtime remains highly competitive, as most of their corresponding points cluster around the $1\times$ line. Overall, this demonstrates that EDQC strikes an excellent balance between achieving state-of-the-art solution quality and maintaining high computational efficiency.
4.5 Statistical Analysis
@@
Figure 3: Critical Difference diagrams for algorithm ranks on all 75 datasets (significance level 0.05). A horizontal bar connects algorithms with no statistically significant difference.
To assess statistical significance, we applied the Friedman test friedman1937use on the algorithm rankings across all 75 datasets. The null hypothesis that all algorithms perform equally was rejected in both settings ($p<0.05$), allowing us to proceed with the post-hoc Nemenyi test demvsar2006statistical to identify pairwise performance differences.
In summary, the statistical tests provide conclusive evidence for the superiority of EDQC. In both experimental settings, EDQC not only achieves the best average rank but also demonstrates a statistically significant performance advantage over all compared state-of-the-art baselines. These results provide strong statistical evidence that EDQC represents a reliably superior approach for this task.
4.6 Analysis of EDQC Components
@@
The results, summarized in Table 3, provide conclusive evidence that each component is essential for EDQC’s superior performance. The most dramatic performance degradation occurred with EDQC-5, which replaced our adaptive diffusion with a naive random process, and EDQC-1, which substituted conductance-based identification with a max-density heuristic. The full EDQC algorithm proved superior in 106 and 69 instances, respectively. These results underscore that our principled diffusion and extraction mechanisms form the foundational pillars of the algorithm’s success. The importance of the energy-guided ranking was also clearly demonstrated; disrupting the descending energy order with a random (EDQC-2) or ascending (EDQC-3) ranking severely hampered performance, leading to 56 and 70 wins for the full EDQC. Finally, disabling the refinement phase (EDQC-4) also resulted in a noticeable performance drop in 16 instances, confirming that this final step is vital for optimizing the solution size.
@@
Table 3: Comparison of the full EDQC with its five simplified variants across 112 experimental settings. #bet and #wor represent the number of instances where EDQC achieves better and worse quasi-clique sizes, respectively.
@@
Figure 4: Subgraph density vs. total retained energy on eight representative datasets. Each blue dot represents one of 1,000 randomly sampled subgraphs of a fixed size; the red marks the quasi-clique of the same size found by EDQC.
4.7 Empirical Validation of the Energy-Density Premise

### v3

@@ 3 $P\leftarrow$ vertices of AA sorted by decreasing $f$;
7 $S\leftarrow\{v_{1},\dots,v_{i^{*}}\}$;
8
9while $|S|>1$ and $\eta(G[S])<\gamma$ do
@@ 23        $\textit{changed}\leftarrow\textbf{true}$;
First, the algorithm collects active vertices $A$ and sorts them by decreasing energy to form an ordered list $P$ (lines 1–2). If $|P|<2$, it returns $\emptyset$ (line 3). This ranking is important, as it places vertices that are more likely to belong to a cohesive region near the front of the sequence.
Next, to obtain an initial candidate set, the algorithm performs a sweep cut (Andersen et al., 2006) on the energy-ranked list $P$ (lines 4–5). Formally, let $P=(v_{1},\dots,v_{k})$. The procedure evaluates the conductance of each prefix $S_{i}=\{v_{1},\dots,v_{i}\}$ for $1\leq i\leq k-1$, and selects the minimum-conductance prefix as the initial set $S$. Note that we evaluate prefixes up to k1k-1 to avoid the degenerate case $S=V$ (or $S=P$ here), for which $\Phi(S)$ is undefined as the denominator becomes zero. In this way, the diffusion ranking is converted into a vertex set whose boundary is identified by the conductance metric. Since $P$ is induced by diffusion scores, low-conductance prefixes are natural candidates for subsequent refinement.
Finally, the resulting initial set $S$ undergoes a refinement procedure (lines 6–18). The refinement begins with a pruning loop (lines 6–8). If the pruning reduces $S$ to a size smaller than $2$, the procedure returns an empty set (line 9), since a $\gamma$-quasi-clique requires at least two vertices. If the initial set $S$ does not meet the density threshold, this loop iteratively removes the vertex with the minimum internal degree $d_{S}(u)$ until the set $S$ induces a valid $\gamma$-quasi-clique. Following this, an expansion phase seeks to enlarge the set while preserving feasibility (lines 10–18). In each pass, the remaining candidates $R=P\setminus S$ are re-sorted based on their connections to the current set $S$ (line 13). The algorithm then iterates through $R$ and greedily adds the first vertex $v$ that maintains the density constraint (lines 15–16). Upon a successful insertion, the current iteration over $R$ is terminated, since the connectivity ordering depends on the updated set $S$. The process repeats until no further insertion is possible without violating $\gamma$.
@@
Let $k=|A|$. ExtractQuasiClique runs in $O(k^{2}+\mathrm{vol}(A))$ time and uses $O(k)$ space.
@@
Sorting the $k$ active vertices takes $O(k\log k)$. The sweep cut linearly scans the sorted list; computing the conductance of all prefixes via incremental updates takes $O(\mathrm{vol}(A))$ time. The refinement stage entails a greedy optimization sequence. In the worst case, it performs $O(k)$ vertex additions or removals. By maintaining internal degrees incrementally, each update requires $O(k)$ operations, yielding an $O(k^{2})$ bound for refinement. Since $k\log k\in O(k^{2})$, the total time complexity is $O(k^{2}+\mathrm{vol}(A))$. The space usage is dominated by the ranked list and auxiliary structures for vertex membership and internal degrees, which use $O(k)$ space. ∎
@@
Fix a source vertex $v$. Let $k=|A|$. One iteration of EDQC with source $v$ runs in $O\!\big(T(k+\mathrm{vol}(A))+k^{2}\big)$ time and uses $O(n+m)$ space. The total time complexity for processing $B$ sources is the summation of this bound over the $B$ iterations. In the degenerate worst case where $B=n$ and $k=\Theta(n)$ in every iteration, this bound recovers $O(n^{3})$.
@@
Consider a fixed source $v$. The diffusion stage runs for $T$ rounds. In our implementation, the active set is maintained incrementally, so each round updates only active vertices and scans their incident edges, incurring $O\!\left(T(k+\mathrm{vol}(A))\right)$ time. The extraction stage sorts the $k$ active vertices in $O(k\log k)$. The subsequent sweep cut requires $O(\mathrm{vol}(A))$ time via incremental updates; since $\mathrm{vol}(A)\leq T(k+\mathrm{vol}(A))$ for $T\geq 1$, this cost is subsumed by the diffusion stage. The refinement stage performs a greedy optimization sequence; with incremental degree maintenance, its worst-case cost is bounded by $O(k^{2})$. Therefore, the total time complexity of one iteration is $O\!\big(T(k+\mathrm{vol}(A))+k^{2}\big)$; in the degenerate worst case where $k=\Theta(n)$ and $B=n$, the total time over all sources becomes $O(n^{3})$. The space usage is $O(n+m)$, dominated by storing $G$ and two reusable energy functions $f$ and $f^{\prime}$. ∎
4. Experimental Evaluation
We conducted a comprehensive experimental evaluation of EDQC. We report solution quality, stability, and runtime under instance-specific density thresholds, together with statistical tests, ablations, and additional analyses on robustness across $\gamma$ and the energy–density premise.
4.1. Experimental Setup
All algorithms were implemented in C++ and compiled with g++ 9.4.0 using the “-O3” optimization flag. Experiments were run on a machine with an AMD EPYC 7H12 CPU (1.5 GHz base frequency), 128 GB of RAM, and Ubuntu 20.04. Baseline parameters followed the original papers. For EDQC, we fixed $T=2$ and $\theta=10^{-4}$ (chosen based on the sensitivity study in Section 4.10) across all datasets.
Due to inherent randomness, all non-deterministic methods (all except NBSim) were executed 10 times with random seeds from 1 to 10; NBSim, being deterministic, was run once. For the non-deterministic methods, we report their average quasi-clique size and standard deviation (STD) over the 10 runs, presented in the format “average size $\pm$ STD”, where the standard deviation is reported only if it is non-zero. All runs were subject to a 60-second timeout. If an algorithm failed to find any valid $\gamma$-quasi-clique in all 10 runs for an instance, we mark its result as “N/A”. The 60-second timeout reflects common practice in recent time-bounded heuristic studies (Lin et al., 2017; Zheng et al., 2023). We report wall-clock time excluding file I/O, and run all methods in a single thread.
4.2. Baseline Methods
@@
  • NBSim/FastNBSim (Pang et al., 2024): Similarity-based heuristics that expand subgraphs based on neighborhood overlap. FastNBSim accelerates NBSim via MinHash-based approximation.
  • NuQClqF1/NuQClqR1 (Liu et al., 2024): Advanced variants of NuQClq (Chen et al., 2021). These algorithms employ information feedback models that integrate historical solutions using fitness-weighted scores to guide local search.
4.3. Datasets
We evaluated all algorithms on 75 real-world datasets from SNAP (Leskovec and Krevl, 2014) and the Network Repository (Rossi and Ahmed, 2015). The collection covers diverse domains, including web, social, and communication graphs. The datasets span a wide spectrum of scales, with vertex counts ($n$) ranging from 62 to over 16.7 million, and edge densities spanning from $8.41\times 10^{-2}$ down to as low as $1.87\times 10^{-7}$.
4.4. Main Results
This section reports the main results on solution quality, stability, and runtime under instance-specific density thresholds induced by NBSim and FastNBSim. A key challenge in this comparison is that NBSim and FastNBSim do not take a density threshold $\gamma$ as input, whereas EDQC, NuQClqF1, and NuQClqR1 do. To enable a comparable evaluation across these methods, we adopt two evaluation settings, following the protocol of Pang et al. (Pang et al., 2024).
Table 1. Quasi-clique sizes with the threshold $\gamma$ set by NBSim’s output.
@@ 4.4. Main Results
In the first setting, NBSim is run once per instance, and the density of its output quasi-clique is used as the target threshold $\gamma$ for EDQC, NuQClqF1, and NuQClqR1. In the second setting, FastNBSim is run 10 times following the protocol in Section 4.1; the density threshold $\gamma$ is then set to the maximum density achieved over these runs, which reflects the strongest cohesion level FastNBSim can reach on the instance, and EDQC, NuQClqF1, and NuQClqR1 are evaluated under this common threshold. The detailed quasi-clique sizes in these two settings are reported in Tables 1 and 2; to save space, instances where all algorithms obtain quasi-cliques of identical average size are omitted.
Table 2. Quasi-clique sizes with $\gamma$ set by FastNBSim’s max density over 10 runs. For FastNBSim, stats are on successful runs only (rate shown in parentheses).
@@
Beyond average size, EDQC demonstrates exceptional stability. It consistently reports a standard deviation close to zero (e.g., $0.32$ on the large eu-2005 graph under the induced $\gamma$), even on complex instances. This is in stark contrast to the NuQClq variants, which frequently exhibit high variance on the same instances (e.g., a STD of $115.84$ for NuQClqF1 on eu-2005). Furthermore, the results highlight the unreliability of the similarity-based methods. FastNBSim, for instance, often fails to achieve its own best-found density (succeeding only 1/10 times on many graphs). This lack of explicit density control makes them less suitable for applications requiring highly cohesive structures, a task where EDQC’s reliability is a distinct advantage.
@@
The efficiency of EDQC is compared against baselines in Figure 2, where, following prior work (Sun et al., 2025), instances where the runtime of both EDQC and the corresponding baseline is below 0.1 seconds are excluded. On this log-log plot, points falling below the $1\times$, $10\times$, and $100\times$ lines indicate that EDQC is faster than the baseline by the corresponding factor or more. The results indicate that EDQC is substantially more efficient than the local search methods, NuQClqF1 and NuQClqR1, with many points falling below the $10\times$ line, signifying a speedup of over an order of magnitude. In comparison to the similarity-based heuristics (NBSim and FastNBSim), EDQC’s runtime remains competitive, as most of their corresponding points cluster around the $1\times$ line. Overall, this demonstrates that EDQC strikes a competitive balance between achieving state-of-the-art solution quality and maintaining high computational efficiency.
Figure 2. Runtime comparison of EDQC against baselines on a log-log scale.
4.5. Statistical Analysis
To assess statistical significance, we applied the Friedman test (Friedman, 1937) on the algorithm rankings across all 75 datasets. The null hypothesis that all algorithms perform equally was rejected in both settings ($p<0.05$), allowing us to proceed with the post-hoc Nemenyi test (Demšar, 2006) to identify pairwise performance differences. On each dataset, we rank algorithms by the reported quasi-clique size under the same induced $\gamma$ (for randomized methods, the mean size over 10 runs). Ties receive averaged ranks. When a method is marked as “N/A” on a dataset, we treat it as the worst performance for ranking on that dataset.
@@ 4.5. Statistical Analysis
Figure 3. Critical Difference diagrams for algorithm ranks on 75 datasets (significance level 0.05). A horizontal bar connects algorithms with no statistically significant difference.
In summary, the statistical tests provide strong evidence that EDQC achieves better average ranks than the compared baselines under both experimental settings at the 0.05 significance level.
4.6. Analysis of EDQC Components
@@
The results, summarized in Table 3, provide strong evidence that each component contributes to the overall performance of EDQC. The most dramatic performance degradation occurred with EDQC-5, which replaced our adaptive diffusion with a naive random process, and EDQC-1, which substituted conductance-based identification with a max-density heuristic. The full EDQC algorithm proved superior in 106 and 69 instances, respectively. These results underscore that our principled diffusion and extraction mechanisms form the foundational pillars of the algorithm’s success. The importance of the energy-guided ranking was also clearly demonstrated; disrupting the descending energy order with a random (EDQC-2) or ascending (EDQC-3) ranking severely hampered performance, leading to 56 and 70 wins for the full EDQC. Finally, disabling the refinement phase (EDQC-4) also resulted in a noticeable performance drop in 16 instances, confirming that this final step is vital for optimizing the solution size.
@@
Table 3. Comparison of the full EDQC with its five simplified variants across 112 experimental settings. #bet and #wor represent the number of instances where EDQC achieves better and worse quasi-clique sizes, respectively.
@@
4.7. Results under Fixed Density Thresholds
Figure 4. Fixed-$\gamma$ comparison of EDQC, NuQClqF1, and NuQClqR1 on 75 graphs for $\gamma\in\{0.7,0.8,0.9,0.95,0.99\}$. Left: normalized solution quality (1.0 = best average quasi-clique size at that $\gamma$). Right: normalized instability (0 = lowest, 1 = highest average standard deviation at that $\gamma$).
Beyond the instance-specific thresholds induced by NBSim and FastNBSim, we also examine how algorithms that accept an explicit density parameter behave across a shared set of fixed thresholds. Specifically, we compare EDQC, NuQClqF1, and NuQClqR1 on all 75 datasets under $\gamma\in\{0.7,0.8,0.9,0.95,0.99\}$. We exclude NBSim and FastNBSim here since they do not take $\gamma$ as input.
For each graph and each $\gamma\in\{0.7,0.8,0.9,0.95,0.99\}$, the three algorithms are run under the same protocol as in Section 4.1. For each dataset and each (algorithm, $\gamma$), we compute the mean and standard deviation (STD) of the quasi-clique size over 10 runs, and then average these quantities over the 75 datasets. If an algorithm fails to find any valid $\gamma$-quasi-clique on a dataset, we treat it as the worst performance for aggregation.
The left panel of Figure 4 shows a radar plot of the normalized solution quality, where for each $\gamma$ the best average size among the three algorithms is scaled to $1.0$ (computed per $\gamma$ across the three algorithms). The right panel reports the normalized instability, where for each $\gamma$ the lowest average STD is mapped to 0 and the highest to $1$ (again per $\gamma$). We use a radar plot to visualize the performance profile across multiple $\gamma$ values.
Figure 5. Subgraph density vs. total retained energy on eight representative datasets. Each blue dot represents one of 1,000 randomly sampled subgraphs of a fixed size; the red marks the quasi-clique of the same size found by EDQC.
The plots reveal a consistent pattern. On the solution-quality radar (left), EDQC stays close to $1.0$ across all density thresholds, indicating that it attains the best (or tied-best) average size at each $\gamma$. Moreover, the gap in normalized size tends to widen as $\gamma$ increases from $0.7$ to $0.99$, suggesting that EDQC is particularly strong when stricter cohesion is required. On the instability radar (right), EDQC remains near 0 across all densities, reflecting consistently low variability, while NuQClqF1 shows higher normalized instability and NuQClqR1 lies in between. These fixed-threshold results provide additional evidence that EDQC delivers both high solution quality and low run-to-run variability over a broad range of density requirements.
4.8. Structural characterization on web-Google
We further examine the structural properties of the $\gamma$-quasi-cliques returned by EDQC and NuQClqF1/R1 on the large web graph web-Google at $\gamma=0.9$. Table 4 reports the mean and standard deviation over 10 runs. EDQC finds substantially larger groups ($60.0\pm 0.0$ nodes vs. $43.8\pm 8.4$ and $41.2\pm 6.5$) while maintaining similarly high internal density. At the same time, its conductance is much lower and far more stable ($0.101\pm 0.003$ vs. $0.570\pm 0.305$ and $0.669\pm 0.235$), indicating that EDQC consistently identifies larger and better isolated cohesive groups. The clustering coefficients are very high and comparable ($\approx 0.92$$0.93$), suggesting similar internal triangle structure across methods.
Table 4. Structural properties of $\gamma$-quasi-cliques on web-Google at $\gamma=0.9$. Mean $\pm$ standard deviation over 10 runs.
Method $|S|$ Density $\Phi(S)$ Clust. coeff.
EDQC $60.0$ $0.904\pm 0.002$ $0.101\pm 0.003$ $0.921\pm 0.001$
NuQClqF1 $43.8\pm 8.4$ $0.911\pm 0.010$ $0.570\pm 0.305$ $0.934\pm 0.015$
NuQClqR1 $41.2\pm 6.5$ $0.907\pm 0.004$ $0.669\pm 0.235$ $0.929\pm 0.015$
4.9. Empirical Validation of the Energy-Density Premise